OpenStudy (anonymous):

Sq rt 1-cos^2 x/ 1+cos^2 x

OpenStudy (anonymous):

\[\sqrt{\frac{ 1-\cos ^{2}x }{ 1+\cos ^{2}x }}\]

OpenStudy (perl):

whats the question

OpenStudy (anonymous):

is it not cos2x ? (in that case the answer is tan x)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):